A 0.18 kg billiard ball moving to the right at 1.2 m/s has a head-on elastic collision with another ball of an equal mass moving to the left at 0.85 m/s.

The first ball moves to the left at 0.85 m/s after the collision.

Find the velocity of the second ball after the collision. Show your work in detail.


1
Expert's answer
2020-11-23T10:31:56-0500

Let's choose the right as the positive direction. Then, we can find the velocity of the second ball after the collision from the Law of Conservation of Momentum:


"m_1v_{1i}-m_2v_{2i}=-m_1v_{1f}+m_2v_{2f},"

here, "m_1=m_2=0.18\ kg" are the masses of the first and second balls, respectively, "v_{1i}=1.2\ \dfrac{m}{s}" is the velocity of the first ball before the collision, "v_{2i}=0.85\ \dfrac{m}{s}" is the velocity of the second ball before the collision, "v_{1f}=0.85\ \dfrac{m}{s}" is the velocity of the first ball after the collision and "v_{2f}" is the velocity of the second ball after the collision.

Then, from this equation we can calculate "v_{2f}":


"v_{2f}=\dfrac{m_1v_{1i}-m_2v_{2i}+m_1v_{1f}}{m_2},""v_{2f}=\dfrac{0.18\ kg\cdot 1.2\ \dfrac{m}{s}-0.18\ kg\cdot 0.85\ \dfrac{m}{s}+0.18\ kg\cdot 0.85\ \dfrac{m}{s}}{0.18\ kg}=1.2\ \dfrac{m}{s}."


The sign plus means that the second ball moves to the right after the collision.

Answer:

"v_{2f}=1.2\ \dfrac{m}{s}," to the right.